# Displacive Jahn–Teller transition in NaNiO2

### v2

@@ Methods
X-ray absorption spectroscopy (XAS) measurements were performed on the BM23 instrument 105 of the European Synchrotron Radiation Facility (ESRF), France. Powdered NaNiO2 was mixed homogeneously with a powdered boron nitride binder ($\sim$16 mg:$\sim$70 mg NaNiO2:BN) and pressed into a 13 mm pellet in a dry, oxygen-free nitrogen environment. The pellet was cut and placed in a sealed sample holder under flowing helium. Temperature control was achieved using a resistive heater, and at each temperature the sample was left for 6 minutes to thermally equilibriate. At each temperature the sample was measured ten times to ensure reproducibility, and an average spectrum produced. X-ray absorption was measured in the vicinity of the Ni K edge ($\sim 8.3$ keV).
X-ray absorption was calculated using the Beer-Lambert law 106 comparing X-ray intensity in ionisation chambers before and after transmission through the sample, and a third ionisation chamber measured the absorption through Ni foil as a reference. Data normalisation was performed using the Python-based 103 Larch 107 package. For the normalisation of the data, a pre-edge range between -350 eV and -45 eV from the edge centre was fit with a straight line, and the post-edge region between 200 eV and 1250 eV was fit with a quadratic polynomial. Data were transformed from $E$-space to $k$-space via the transform $k=\sqrt{\frac{2m_{e}}{\hbar^{2}}\left(\hbar\omega-E_{0}\right)}$, where $m_{e}$ is the mass of the electron, $E_{0}$ is the edge energy, and $\omega$ is the frequency of the measured photon. Finally, the EXAFS function $\chi(k)$ was obtained by fitting and subtracting a spline to the EXAFS data in $k$-space, where $R_{\mathrm{bkg}}$ was set to 1 Å and the spline was fit within a $k$-range of 0 Å-1 to 18.25 Å-1, with a $k^{2}$-weighting. For the Fourier transform of EXAFS from $k$-space to $R$-space, a $k$-range of 2.5 Å-1 to 15 Å-1 was used, and to minimise Fourier ripples from the cutoffs in $k$-space, a Kaiser-Bessel window with end width $dk=7$-1 was applied to the data in $k$-space before the transformation.
Calculated EXAFS data were obtained by multiple-scattering path expansion using FEFF 48, as implemented in Larch in Python. In this model, data is fit using the EXAFS equation over each set of equivalent paths $j$, defined as follows:
$\chi(k)=S_{0}^{2}\sum_{j}\frac{N_{j}|f_{j}(k)|}{kR_{j}^{2}}\sin{\left[2kR_{j}+\phi_{j}(k)\right]}\times\exp{\left[-\frac{2R_{j}}{\lambda_{j}(k)}\right]}\exp(-\sigma_{j}^{2}k^{2})$ (3)
@@
AIMD simulations were performed according to the Generalized Gradient Approximation (GGA) 108 and the projector augmented wave method (PAW) 109, as implemented in the Vienna Ab Initio Simulation Package (VASP) 110, 111. The plane-wave energy cutoff was set to 500 eV. Supercells are expansions of the monoclinic ($3\times 3\times 3$; 216 ions) and rhombohedral ($4\times 4\times 1$; 192 ions) unit cells, respectively. Simulations were performed at the $\Gamma$ point and checked against calculations with a $2\times 2\times 2$ Monkhorst-Pack 112 k-point grid for consistency. The convergence criteria for the electronic and ionic relaxations were set to $10^{-6}$ eV and $5\cdot 10^{-3}$ eV/Å, respectively.
For Ni, the $4s^{2}3d^{8}$ electrons were treated as valence electrons. To account for the strongly correlated d electrons, a rotationally invariant Hubbard U parameter 113 of $U_{\mathrm{eff}}=6$ eV was selected, which was used successfully in previous studies of layered oxide cathodes, both for 0 K DFT calculations and finite temperature AIMD simulations 114, 115, 17. For oxygen, the $2s^{2}2p^{4}$ electrons were considered as valence electrons.
AIMD simulations were performed for the isothermal-isobaric ensemble (NpT, constant pressure, particle number, and temperature) at zero pressure, allowing the cell shape and volume to change. A Langevin thermostat was used with friction coefficients set to zero to minimise impact on the lattice vibrations. Trajectories were run for 3 ps at timesteps of 1 fs. Following an equilibration period of 1 ps, snapshots were sampled every 2-5 fs for the analysis of the distortions.
@@ Acknowledgements
Crystal structure figures were prepared using Vesta-3 116.
@@
Data availability
Data from the BM23 instrument at the European Synchrotron Radiation Facility is available at doi:10.15151/ESRF-ES-962076745 117.
All other data is available in the University of Cambridge repository at doi:[XXXXXXX] .
Supplementary Information
A document containing Supplementary Information to this article is available [Link to PDF here].

### v4

@@ Methods
X-ray absorption spectroscopy (XAS) measurements were performed on the BM23 instrument 106 of the European Synchrotron Radiation Facility (ESRF), France. Powdered NaNiO2 was mixed homogeneously with a powdered boron nitride binder ($\sim$16 mg:$\sim$70 mg NaNiO2:BN) and pressed into a 13 mm pellet in a dry, oxygen-free nitrogen environment. The pellet was cut and placed in a sealed sample holder under flowing helium. Temperature control was achieved using a resistive heater, and at each temperature the sample was left for 6 minutes to thermally equilibriate. At each temperature the sample was measured ten times to ensure reproducibility, and an average spectrum produced. X-ray absorption was measured in the vicinity of the Ni K edge ($\sim 8.3$ keV).
X-ray absorption was calculated using the Beer-Lambert law 107 comparing X-ray intensity in ionisation chambers before and after transmission through the sample, and a third ionisation chamber measured the absorption through Ni foil as a reference. Data normalisation was performed using the Python-based 104 Larch 108 package. For the normalisation of the data, a pre-edge range between -350 eV and -45 eV from the edge centre was fit with a straight line, and the post-edge region between 200 eV and 1250 eV was fit with a quadratic polynomial. Data were transformed from $E$-space to $k$-space via the transform $k=\sqrt{\frac{2m_{e}}{\hbar^{2}}\left(\hbar\omega-E_{0}\right)}$, where $m_{e}$ is the mass of the electron, $E_{0}$ is the edge energy, and $\omega$ is the frequency of the measured photon. Finally, the EXAFS function $\chi(k)$ was obtained by fitting and subtracting a spline to the EXAFS data in $k$-space, where $R_{\mathrm{bkg}}$ was set to 1 Å and the spline was fit within a $k$-range of 0 Å-1 to 18.25 Å-1, with a $k^{2}$-weighting. For the Fourier transform of EXAFS from $k$-space to $R$-space, a $k$-range of 2.5 Å-1 to 15 Å-1 was used, and to minimise Fourier ripples from the cutoffs in $k$-space, a Kaiser-Bessel window with end width $dk=7$-1 was applied to the data in $k$-space before the transformation.
Calculated EXAFS data were obtained by multiple-scattering path expansion using FEFF 49, as implemented in Larch in Python. In this model, data is fit using the EXAFS equation over each set of equivalent paths $j$, defined as follows:
$\chi(k)=S_{0}^{2}\sum_{j}\frac{N_{j}|f_{j}(k)|}{kR_{j}^{2}}\sin{\left[2kR_{j}+\phi_{j}(k)\right]}\\ \times\exp{\left[-\frac{2R_{j}}{\lambda_{j}(k)}\right]}\exp(-\sigma_{j}^{2}k^{2})$ (3)
@@
AIMD simulations were performed according to the Generalized Gradient Approximation (GGA) 109 and the projector augmented wave method (PAW) 110, as implemented in the Vienna Ab Initio Simulation Package (VASP) 111, 112. The plane-wave energy cutoff was set to 500 eV. Supercells are expansions of the monoclinic ($3\times 3\times 3$; 216 ions) and rhombohedral ($4\times 4\times 1$; 192 ions) unit cells, respectively. Simulations were performed at the $\Gamma$ point and checked against calculations with a $2\times 2\times 2$ Monkhorst-Pack 113 k-point grid for consistency. The convergence criteria for the electronic and ionic relaxations were set to $10^{-6}$ eV and $5\cdot 10^{-3}$ eV/Å, respectively.
For Ni, the $4s^{2}3d^{8}$ electrons were treated as valence electrons. To account for the strongly correlated d electrons, a rotationally invariant Hubbard U parameter 114 of $U_{\mathrm{eff}}=6$ eV was selected, which was used successfully in previous studies of layered oxide cathodes, both for 0 K DFT calculations and finite temperature AIMD simulations 115, 116, 17. For oxygen, the $2s^{2}2p^{4}$ electrons were considered as valence electrons.
AIMD simulations were performed for the isothermal-isobaric ensemble (NpT, constant pressure, particle number, and temperature) at zero pressure, allowing the cell shape and volume to change. A Langevin thermostat was used with friction coefficients set to zero to minimise impact on the lattice vibrations. Trajectories were run for 3 ps at timesteps of 1 fs. Following an equilibration period of 1 ps, snapshots were sampled every 2-5 fs, using Ovito 117, for the analysis of the distortions.
@@
Data availability
Data from the BM23 instrument at the European Synchrotron Radiation Facility is available at doi:10.15151/ESRF-ES-962076745 118. All other data is available in the University of Cambridge repository at doi.org/10.17863/CAM.112349 119.
Supplementary Information
A document containing Supplementary Information to this article is available. The Supplementary Information contains: A description of the Van Vleck modes. Further analysis of supercell outputs of big box PDF analysis, and results from various different starting configurations. Expanded AIMD results. NMR data. Tabulated results of all analysis techniques.
@@ Acknowledgements
Crystal structure figures were prepared using Vesta-3 120 and graphs were prepared using MatPlotLib 121.